OpenStudy (anonymous):

1. Which ordered pair is a solution of the equation y = x -4? a. (2,6) b. (6,2) c. (-2,6) d. (-6,2) 2. Which ordered pair is a solution to the equation y =5x? a. (-2,10) b. (-5,25) c. (-3,15) d. (-2,-10)

OpenStudy (shamil98):

y = x - 4 y - x = -4 2 - 6 = -4 (-6,2)

OpenStudy (shamil98):

y = 5x -10 = 5(-2) (-2,-10) just plug in the choices you are given and see what works

OpenStudy (anonymous):

Thank you!! :)